Find the approximately the difference between the areas of two spheres whose radii are 4 feet and 4.05 feet.


1
Expert's answer
2021-01-19T04:36:04-0500

We have "r_1=4ft" and "r_2=4.05ft"

And total surface area of a sphere is

"S=4\pi r^2"

,where "r" is the radius of the sphere.


So, difference between the two surface areas:-


"S_2-S_1\\
=4\pi r_2^2-4\pi r_1^2\\
=4\pi(r_2^2-r_1^2)\\
=4\pi((4.05)^2-(4)^2)\\
=4\pi(16.4025-16)\\
=4\pi(0.4025)\\
=4\cdot3.14\cdot0.4025\\
=5.0554\>sq\>ft"

So, the difference between the surface areas is "5.0554" sq ft.